OpenStudy (anonymous):

the area of a soccer field is 5500 square yards.the width of the field is 60 yards less than the length.find the length,in yards,of the field hockey field.

OpenStudy (pfenn1):

If x is the length of the field, the width is given by x-60. The area of the soccer field would be given by the width multiplied by the length.

OpenStudy (anonymous):

so i divied

OpenStudy (pfenn1):

divied?

OpenStudy (anonymous):

i Met Multiply

OpenStudy (pfenn1):

So are you all set?

OpenStudy (anonymous):

no can u show an example on the drawing board

OpenStudy (pfenn1):

Your equation will be \[x(x-60)=5500\]\[ x^2-60x-5500=0\]which is a binomial. Solve for x.

OpenStudy (pfenn1):

Or factor the equation \[x^2-60x-5500=(x-110)(x+50)=0\]